A lease has a term of 8 years with annual payments of $24,100. The asset would cost $118,000 to buy and would be depreciated straight-line to a zero salvage value over 8 years. The actual salvage value is zero. If the firm has a tax rate of 25 percent, what is the incremental cash flow in Year 8 of leasing rather than purchasing? $14,856.50 -$16,375.00 -$25,418.50 $16,375.00 -$21,762.50
Answer #1

Incremental expense in Year 8 of leasing = Amount paid in purchase - Lease Rent

= 24,100- 118000/8

= $9,350

Tax savings = 9,350*25% = $2,337.5

Hence, incremental cash flow = -24,100+2,337.5 = -$21,762.50

Since depreciation is a non cash cost

Hence, the answer is -$21,762.50